OpenStudy (maeganr):

Write an equation in point-slope form of the line through the given points. Then write the equation in slope-intercept form. #11. (4,0), (-2,1) #12. (-3,-2), (5,3) #13. (-5,1), (3,4) Thats all they give you. And YES I do know the Formulas: Point-slope: y-y1=m(x-x1) Slope-intercept: y=mx+b Slope: m=y2-y1/x2-x1

OpenStudy (chihiroasleaf):

first.., find the slope using the third formula that you write.., then use point-slope formula to find the equation..

OpenStudy (anonymous):

m=y2-y1/x2-x1 (4, 0) (-2, 1) m=1-0/-2-4 m=1/-6 m=y2-y1/x2-x1 (-3, -2), (5, 3) m=3-(-2)/5-(-3) m=5/8 m=y2-y1/x2-x1 (-5, 1) (3, 4) m=4-1/3-(-5) m=3/8

OpenStudy (maeganr):

SLOPE= -0.16

OpenStudy (maeganr):

@Butterfly16 Thats what I've done before, but then people come along, and tell me I'm wrong! ):

OpenStudy (maeganr):

And why are you finding the slope 3 times?

OpenStudy (chihiroasleaf):

because you have three questions and each question have equation... :)

OpenStudy (maeganr):

oh, well I'm on #11.

OpenStudy (anonymous):

y-y1=m(x-x1) (4,0), (-2,1) y-0=(1/-6)(x-4) y-0= -0.16x+0.7 y=-0.16x+0.7 Yeah, I just finding the slope for all of them. :o)

OpenStudy (maeganr):

Im so freaking confused.

OpenStudy (chihiroasleaf):

now.., using the similar way.., you can solve questions #12 and #13 can you? :)

OpenStudy (anonymous):

y1 and x1 just stand for the points. So (4, 0) is the point, so y1 is 0 and x1 is 4. That would the equation for the point-slope form for that set of points.

OpenStudy (maeganr):

but you have to write it in point-slope form?!

OpenStudy (anonymous):

Then, I guess you can just plug in the numbers into the point-slope form so it'd just be: y-0=(1/-6)(x-4)

OpenStudy (maeganr):

but don't you have to do 2 equations because of (4,0) and (-2,1)

OpenStudy (chihiroasleaf):

@Butterfly16 has write it in slope-intercept form slope-intercept form is y = mx+ c , No.., points (4,0) and (-2,1) are on the same line.., so just choose one of them..

OpenStudy (maeganr):



OpenStudy (maeganr):

because like that?

OpenStudy (anonymous):

Yip, the line would look like that. :o)

OpenStudy (maeganr):



OpenStudy (chihiroasleaf):

yup.., it's for the second line... :)

OpenStudy (anonymous):

Awesome job! :o)

OpenStudy (maeganr):

y-y1=m(x-x1) y--2=1/2(x--3) help

OpenStudy (anonymous):

The points for number 12?

OpenStudy (maeganr):

@erdog82 Kinda.

OpenStudy (maeganr):

#12. i got that far but I'm stuck now.

OpenStudy (anonymous):

y-y1=m(x-x1) (-3, -2) (5, 3) y-(-2)=(5/8)(x-(-3)) y+2=0.625x-1.875 y=0.625x-3.875

OpenStudy (maeganr):

but y=mx+b

OpenStudy (anonymous):

Yup, so for #12 it'd be y=0.625x-3.875